1.SIMPLIFY AND WRITE WITH POSITIVE EXPONENTS 9X(X^-3)/18X^15
I DID THIS AND NOW IM STUCK:9X(1/X^-3) THEN I GOT 9X/18X^15X^-3 = 9X/18X^12 = X/2X^12 NOW IM STUCK ...HELP?
2.WRITE WITHOUT NEGATIVE EXPONENTS=(3X^-4Y^2)^-3 = (1/3X^-4Y^2)^-3 = 3*3*3 X^-4+3 Y^2+3 = 1/27X^-1Y^5 I NEED HELP IM ALSO STUCK.

Answer by stanbon(75887)

1.SIMPLIFY AND WRITE WITH POSITIVE EXPONENTS 9X(X^-3)/18X^15
= (9/18)(x*x^-3/x^15)
---
= (1/2)(x^-2/x^15)
---
= (1/2)(1/x^17)
---
= 1/(2x^17)
===============================================

2.WRITE WITHOUT NEGATIVE EXPONENTS=(3X^-4Y^2)^-3
= 3^-3 * x^12 * y^-6
---
= (1/27)(x^12/y^6)
---
= x^2/(27y^6)
================================================
Cheers,
Stan H.